OpenStudy (anonymous):

For what value of x is (x, –3) a solution for 4x – 3y = 21?

OpenStudy (anonymous):

sorry x = 21-9 x=12

OpenStudy (anonymous):

schute i had answered ur question (previous) pls now review on it i have to go pls now

OpenStudy (anonymous):

4x - 3(-3) = 21 4x + 9 = 21 x = 21- 9 x = 12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but give me medal also there

OpenStudy (anonymous):

12 isn't an answer choice

OpenStudy (anonymous):

A11 B–3 C–11 D3

OpenStudy (anonymous):

@schulte pls recheck if the question posted is exactly in the book...

OpenStudy (anonymous):

Ok, i found the mistake i made..... Sorry !! 4x - 3(-3) = 21 4x + 9 = 21 4x = 21- 9 4x = 12 x = 12/4 = 3 so D) 3 is the correct answer sorry one more for the inconvenience and confusion i created.......